OpenStudy (tw101):

Find the slope-intercept form for the equation of the line which passes through the point (−1, 3) and has a slope of −2. y = -2x - 1 y = -2x - 3 y = -2x + 1 y = -2x + 3

OpenStudy (mathmale):

First step: find and type in the "point slope form of the equation of a straight line." Next: Substitute slope value -2 for m in this equation. Substitute the x- and y-coordinates of the given point into this equation. Simplify the result.

OpenStudy (tw101):

is that y-y1=m(x-x1)? @mathmale

OpenStudy (tw101):

so would i put y-3=2(x-(-1))?

OpenStudy (mathmale):

That's great. You have been given the value of the slope, m, as well as the coordinates of one point (x, y). Go ahead and substitute all three values into this equation.

OpenStudy (mathmale):

Yes! except that the slope, m, is -2, not +2.

OpenStudy (tw101):

Oh right! Sorry. One moment:)

OpenStudy (tw101):

Is it y=-2x+1? @mathmale

OpenStudy (mathmale):

Yes. Does this result match any of the answer choices?

OpenStudy (tw101):

Yes!

OpenStudy (tw101):

Thank you very much

OpenStudy (tw101):

Would it be okay if I asked one more question?

OpenStudy (mathmale):

sure, please go ahead

OpenStudy (tw101):

Thanks. Ill attach it.

OpenStudy (tw101):

How do I substitute for point of origin?

OpenStudy (mathmale):

All four of the answer choices are in slope-intercept form. "Intercept" most commonly refers to the y-value at which the line crosses the y-axis. Reading the problem again, what y value do you obtain for the y-intercept / vertical intercept?

OpenStudy (tw101):

2?

OpenStudy (mathmale):

First, would you please write out your conception of the "slope-intercept form?"

OpenStudy (tw101):

y=mx+b

OpenStudy (mathmale):

Right.

OpenStudy (mathmale):

I*f the line passes thru the origin, what is the y-value there?

OpenStudy (mathmale):

Hint: what are the coordinates of the origin?

OpenStudy (tw101):

They're (-3,2)

OpenStudy (tw101):

x=-3 y=2

OpenStudy (mathmale):

that is the other point. The coord. of the origin are (0,0).

OpenStudy (tw101):

Oh right I see

OpenStudy (mathmale):

So, now you have y=mx+b, and know that the line passes thru the origin. b is the y-coordinate of the point at which the line crosses the y-axis. Here, b = ?

OpenStudy (tw101):

So then that means you just have to add the slope into the formula

OpenStudy (tw101):

2

OpenStudy (mathmale):

Yes, you'll need to calculate the slope and insert your result into the formula.

OpenStudy (mathmale):

What happened to the 3?

OpenStudy (tw101):

0

OpenStudy (tw101):

Wait, sorry

OpenStudy (mathmale):

The line passes thru (-3,2) as well as thru (0,0).

OpenStudy (mathmale):

Would this line have a pos. or a neg. slope, and how would you know that?

OpenStudy (tw101):

negative? because the 3 is negative?

OpenStudy (tw101):

Sorry im new at this and just a bit confused

OpenStudy (mathmale):

I'd prefer you actually see the line involved here. It goes thru (-3,2) and (0,0).

OpenStudy (mathmale):



OpenStudy (tw101):

so then would it be -3/2x?

OpenStudy (mathmale):

does the function seem to be increasing or decreasing as x increases?

OpenStudy (tw101):

Decreasing

OpenStudy (mathmale):

Slope = rise / run. Here, the change in y as you progress from (-3,2) to (0,0) is -2 and the change in x is 3. So, your slope is rise / run = -2/3

OpenStudy (mathmale):

Slope does not involve x here; the slope is a constant coefficient in front of the "x". x is your independent variable.

OpenStudy (tw101):

Okay

OpenStudy (mathmale):

So, in this case, the slope is... ? and the y-intercept, b, is ... ? and the correct choice of answer is ... ?

OpenStudy (tw101):

So if it's rise over run -2/3, would y=-2/3x?

OpenStudy (tw101):

If the y-intercept is 0

OpenStudy (mathmale):

Yes, if " b " equals zero, then the entire equation of the line boils down to y=(-2/3) x.

OpenStudy (tw101):

Okay, I definitely understand better now. Thank you very much.

OpenStudy (mathmale):

I'm so glad to have been of help. Good luck to you!